A company borrows $160000, which will be paid back to the lender in one payment at the end of 5 years. The company agrees to pay semi-annually interest payments at the nominal annual rate of 10% compounded semi-annually. At the same time the company sets up a sinking fund in order to repay the loan at the end of 5 years. The sinking fund pays interest at an annual nominal interest rate of 4% compounded semi-annually. Find the total amount of the semi-annually payments, that is, the sum of the interest payment and the sinking fund payment.
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

BC $160,000.00 Loan Amount Term in years Term in years Annual interest rate Semiannual interest rate Semiannual interest paym

Cell reference -

B 160000 Loan Amount Term in years Annual interest rate Semiannual interest rate Semiannual interest payment 0.1 =C4/2 =C2*C5

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.